A series RL circuit with L = 3.80 H and a series R
Answer #1

time constant of RL circuit

tau = L/R

time constant of RC circuit

tau = RC

given that

L/R = RC

R^2 = L/C

R = sqrt (L/C)

R = sqrt (3.8/(3.8*10^-6)) = 1000 ohm = 1 k-ohm

2.

tau = L/R = 3.8/1000 = 3.8*10^-3 sec = 3.8 ms

OR
tau = RC = 3.8*10^-6*1000 = 3.8*10^-3 = 3.8 ms